### v1

# Fait: Fault-Aware Fine-Tuning for Better Code Generation

Abstract.
Modern instruction-tuned large language models (LLMs) have made remarkable progress in code generation. However, these LLMs fine-tuned with standard supervised fine-tuning (SFT) sometimes generate plausible-looking but functionally incorrect code variants. This issue likely stems from the limitation of standard SFT, which treats all tokens equally during optimization and fails to emphasize the error-sensitive segments—specific code differences between correct implementations and similar incorrect variants. To address this problem, we propose Fault-Aware Fine-Tuning (Fait), a novel fine-tuning technique that enhances LLMs’ code generation by (1) extracting multi-granularity (line/token-level) differences between correct and incorrect yet similar implementations to identify error-sensitive segments, and (2) dynamically prioritizing those segments during training via dynamic loss weighting. Through extensive experiments on seven LLMs across three widely-used benchmarks, our method achieves an average relative improvement of 6.9% on pass@1 with just one epoch of training, with some enhanced 6.7B LLMs outperforming closed-source models, e.g., GPT-3.5-Turbo. Furthermore, our fine-tuning technique demonstrates strong generalization with performance improvements ranging from 3.8% to 19.1% across diverse instruction-tuned LLMs, and our ablation studies confirm the contributions of different granularities of differences and loss function components.
1. Introduction
Recently, fine-tuning LLMs using synthetic datasets generated by teacher models has emerged as a popular paradigm for improving code generation capabilities (Luo et al., [n. d.]; Wei et al., 2024; Zheng et al., 2024; Gunasekar et al., 2023; Yu et al., 2024). This paradigm uses stronger teacher models to generate high-quality instruction-response pairs and construct a dataset. These synthetic datasets are then used to fine-tune weaker student models with standard SFT method, which uses instructions to guide LLMs to generate outputs matching reference responses by minimizing cross-entropy loss uniformly across all tokens.
Although these LLMs fine-tuned with standard SFT achieve impressive performance on code generation benchmarks such as HumanEval (Chen et al., 2021), they sometimes generate plausible-looking but incorrect code variants (Huang et al., 2025; Liu et al., 2024a). For example, Llama-3.1-70B-Instruct (Dubey et al., 2024), a model fine-tuned from Llama-3.1-70B using standard SFT, is capable of solving 82.3% of the problems in HumanEval. However, our analysis shows that 34.5% of its failed cases are attributed to deviations in error-prone segments of otherwise correct implementations. As shown in Figure 1, in a vowel identification task, although the LLM successfully implements the core logic of searching for vowels between consonants from the right, it erroneously initializes the starting index of the loop–a mistake in a critical part of the code that renders the program incorrect. We refer to such crucial differences between correct implementations and similar incorrect variants as error-sensitive segments, such as len(word) - 1 and len(word) - 2 in Figure 1. These error-sensitive segments act as critical decision points in code generation, where even slight deviations can determine the correctness of the output.
To address the problem of errors in code generation, we propose Fault-Aware Fine-Tuning (Fait) as a finetuning technique specifically designed to guide LLMs to focus on error-sensitive segments and avoid mistakes in these critical regions, thereby improving the reliability and accuracy of code generation. Implementing this approach involves two main challenges. The first challenge is to identify these error-sensitive segments, as existing instruction-tuning dataset construction methods primarily focus on generating instruction-response pairs without specifically considering these segments (Zhang et al., 2023). To overcome this, we develop a two-phase segment identification component. First, we leverage a teacher model with a carefully designed prompt to generate functionally incorrect yet similar variants of the correct implementations from the existing instruction tuning dataset. Then, we can identify the differences between these correct implementations and their similar yet incorrect variants, which serve as error-sensitive segments. We annotate the tokens in the segments through a multi-granularity method at both line and token levels. Notably, we annotate both the tokens in correct implementations and their corresponding parts in similar yet incorrect variants. The second challenge lies in guiding LLMs to focus on these labeled segments during fine-tuning, as standard SFT treats all tokens with equal importance regardless of their criticality. To address this challenge, we adjust the loss of SFT to prioritize the annotated error-sensitive tokens within correct implementations. Specifically, Fait processes both correct and incorrect implementations to discriminate the error-sensitive segments, and only computes the loss based on correct code implementations. Unlike standard SFT that uniformly weights all tokens during loss computation, we dynamically assign relatively higher weights to those tokens in the correct implementation that correspond to the error-sensitive segments. This methodology enhances LLMs’ capability to discriminate error-sensitive segments when solving programming tasks, thereby increasing the likelihood of generating correct implementation details while suppressing error-prone alternatives.
To implement our method, we construct a refined dataset derived from the original instruction-tuning data. Each data point consists of an instruction, its correct implementation from the original dataset, and an LLM-generated similar incorrect variant. We then develop a multi-granularity error-sensitive segment extraction method and combine it with the refined loss function to enhance LLM’s code generation capabilities.
We validate the effectiveness of the Fait through extensive experiments. Notably, through Fait, the selected LLMs undergo only one epoch of training on their original instruction datasets, yet achieve an average relative improvement of 6.9% on pass@1 (a strict evaluation metric measuring the ratio of first-generated samples that pass all test cases) across three representative code generation benchmarks (Humaneval(+), MBPP(+), and BigCodeBench) (Chen et al., 2021; Austin et al., 2021; Liu et al., 2024b; Zhuo et al., 2024). Among these LLMs, SemCoder-S (Ding et al., 2024) with 6.7B parameters outperforms closed-source models like GPT-3.5-Turbo (OpenAI, 2022) on Humaneval(+) and MBPP(+) benchmarks and MagiCoder-DS with 6.7B parameters outperforms GPT-3.5-Turbo on HumanEval(+). Our method also demonstrates strong generalization capabilities, showing performance improvements ranging from 3.8% to 19.1% across multiple instruction-tuned LLMs, including those trained on closed-source instruction datasets. Moreover, our ablation experiments on Fait confirm the contributions of different granularities of differences in generated code details and loss functions.
We summarize our contributions as follows.
To the best of our knowledge, we are the first to investigate how to enhance LLMs’ understanding of error-sensitive segments by refining the SFT process to improve LLMs’ code generation capabilities.
We propose a novel framework, Fault-Aware Fine-Tuning (Fait), to effectively guide LLMs to focus on error-prone parts in code. This is achieved by (1) extracting multi-granularity code differences (token-/line-level) to identify error-sensitive segments, and (2) refining SFT to dynamically assign higher weights to these parts during the training process.
Through extensive experiments across seven LLMs and three widely-used code generation benchmarks, we demonstrate the effectiveness and generalizability of our approach in effectively boosting LLMs’ code generation performance compared to baseline methods.
2. Approach
Figure 2 illustrates the overview of Fait. This approach takes as input an instruction-tuned LLM and its corresponding instruction-tuning dataset, and outputs an enhanced LLM with improved code generation capabilities that can better discriminate error-sensitive segments. It first augments the dataset by generating similar yet incorrect implementations for each correct response. Then, it identifies error-sensitive segments between the paired implementations and calculates weights for tokens in the correct implementations that differ from the incorrect variants. During fine-tuning, it only computes loss on the correct implementations, with higher weights assigned to tokens in error-sensitive segments, producing an LLM that can better discriminate these error-sensitive segments. The methodology consists of two key components:
Error-Sensitive Segments Identification: This component creates a refined dataset of paired correct and similar yet wrong code samples from the original dataset, and processes code differences at multiple granularities to identify error-sensitive segments.
Dynamic Importance Reweighting: This component strategically reweights token weights in the loss function to prioritize discriminative elements in correct implementations, building upon the identified error-sensitive segments. This dynamic weighting method enhances the LLM’s attention to the key implementation details in correct code, effectively teaching it to distinguish between valid solutions and their similar yet incorrect counterparts, ultimately improving code generation capabilities.
The two components work together to fine-tune LLMs to distinguish between correct implementations and similar yet incorrect alternatives, thereby improving code generation performance. The following subsections detail each component of our methodology.
2.1. Error-Sensitive Segments Identification
The input to this component is the instruction-tuning dataset , where represents the target problem description and denotes the correct implementation. The output is an enhanced dataset with error-sensitive segment information, where represents the corresponding similar but incorrect implementation. To generate incorrect code variants, we utilize a stronger teacher LLM with a carefully designed prompt. The prompt template is shown in Figure 3, which consists of two parts. The first part defines the task for producing incorrect responses corresponding to the target problem and correct answer, explicitly specifying that outputs must be similar to correct solutions, with responses constrained to markdown formatting for consistent post-processing. The second part provides contextual references by presenting the target problem description and the correct solution.
After obtaining pairs of correct and incorrect code, we extract the differences to identify error-sensitive segments and process them at different granularity levels to capture both line-level and token-level information. Specifically, we designate as the pre-change version and as the post-change version.
Line-Level Differences. We align and line-by-line using Python’s difflib library. For each line, the tool assigns a flag indicating whether it should be deleted (-), added (+), or remain unchanged. We extract the lines marked for deletion from and those marked for addition from .
Let and denote the number of code lines in the correct code and incorrect code , respectively. Based on these extracted lines, we construct the line-level boolean mask vectors and for and as follows:
where is the indicator function that outputs 1 if the condition is true and 0 otherwise.
Token-Level Differences. We utilize the Levenshtein distance algorithm (Yujian and Bo, 2007) to identify character-level change information between and . The Levenshtein distance algorithm, also known as the edit distance algorithm, quantifies the minimum number of single-character operations (insertions, deletions, or substitutions) required to transform one string into another. We identify the characters that need to be edited to transform the original string () into the modified version (), and record their positions accordingly. For instance, if a character operation is an insertion, we record its position in , as shown in Figure 2. Given that the LLM’s embedding layer is tightly coupled with the LLM’s tokenizer vocabulary, we map these character-level differences to tokens using the LLM’s tokenizer. When character modifications span multiple tokens, all affected tokens are marked.
Let and denote the number of tokens in and . We construct token-level boolean mask vectors and for and as follows:
Hybrid Level Vectors. To create comprehensive representations of error-sensitive segments, we combine line-level and token-level masks. However, these two types of masks operate at different granularities and cannot be directly combined. We must first align these representations to the same granularity to enable their integration.
We convert line-level masks to token-level granularity, as illustrated in Algorithm 1. Lines 1-2 initialize a token-level mask vector (size T) and the position index. Lines 3-16 implement the core logic - for each code line (1 to L), when the line is masked (marked lines, Lines 5-8), all tokens in that line receive mask value 1; otherwise (unmarked lines, Lines 10-15), tokens receive value 0. For and , we identify the corresponding tokens within that line to get and .
With both masks now represented at the token level, we then use an element-wise addition operation to combine and , as follows:
These hybrid vectors precisely identify error-sensitive segments at multiple granularities, highlighting critical differences between correct and incorrect implementations. Noted that changed tokens must appear in changed lines, our hybrid representation naturally creates a priority system: 1) tokens that are both in changed lines and are themselves changed will have a value of 2 in the hybrid vector; 2) tokens that are only in changed lines but not directly changed will have a value of 1. This provides a more comprehensive view than either granularity alone, with higher values indicating more critical tokens.
2.2. Dynamic Importance Reweighting
With the identified error-sensitive segments, we now refine the SFT process to prioritize these critical differences. Based on the constructed dataset , the standard SFT loss is computed as:
where denotes the number of samples in a batch. Notably, the standard SFT loss function treats all tokens equally.
In contrast, the Fait introduces dynamic token-level weights emphasize error-sensitive segments:
The weight is computed as follows: Given input , outputs and , we first obtain the LLM’s prediction probabilities for both correct and incorrect implementations given the same instruction:
where represents the conditional probability function of the LLM that computes the probability of the next token given the input and previous tokens. We then apply the hybrid-level vectors to isolate probabilities for error-sensitive segments:
Where denotes element-wise multiplication. Inspired by the Bradley–Terry model (Hunter, 2004), a pairwise comparison framework widely used in ranking systems (Baker and Scarf, 2021; Yan, 2016; Menke and Martinez, 2008), we compute dynamic token weights for differentiating tokens in error-sensitive segments:
Where is a hyperparameter controlling the weight range, denotes the mean probability of tokens in error-sensitive segments in , and represents the corresponding value for . This formulation ensures that: (1) When the mean probabilities and are close (indicating the LLM struggles to distinguish the tokens between and ), the weights for differentiating tokens in approach , thereby maximizing emphasis on error-sensitive segments. (2) Conversely, when and diverge significantly (demonstrating the LLM can discriminate the differentiating tokens in and ), the weights diminish toward , reducing emphasis. For tokens shared between and , we assign fixed weights , ensuring the LLM maintains baseline attention to shared elements while prioritizing discriminative features.
This dynamic weighting mechanism guides the LLM to focus on the challenging discriminative aspects of correct implementations, which can improve its code generation capability.
3. Experiments Setup
3.1. Benchmarks
We conduct experiments on three widely used code generation benchmarks to demonstrate the superiority and generality of proposed Fait.
Humaneval (Chen et al., 2021): This benchmark consists of 164 manually crafted programming tasks, created by OpenAI. Each task includes a method signature, a docstring, a method body, and several unit tests. Our work employs both the initial HumanEval and its extended version, HumanEval+(Liu et al., 2024b), which expands the test cases of the original with 80 additional test samples to overcome limitations in test coverage (Liu et al., 2024b).
MBPP (Austin et al., 2021): This benchmark contains 974 Python coding tasks spanning core programming concepts, library utilization capabilities, and more. Our study adopts the extended versions proposed by (Liu et al., 2024b), including MBPP and MBPP+. These collections each contain 378 tasks, with the enhanced version incorporating 35 times the number of test samples.
Bigcodebench (Zhuo et al., 2024): This benchmark presents a rigorous benchmark for code generation, constructed to measure LLMs’ capabilities in utilizing programming tools and the following of complex instructions. It contains 1,140 code-generation problems. In the Complete configuration, each problem provides a function signature, a problem description, and a test suite. A small high-quality subset known as BigCodeBench-Hard contains 148 problems that are more user-centric and challenging. Our study uses both the full set and the hard set, namely, BigCodeBench-Full and BigCodeBench-Hard.
3.2. Metrics
To evaluate code generation performance, we use the Pass@K metric, which is widely used in prior studies (Chen et al., 2021; Mu et al., 2024; Huang et al., 2023). This metric checks whether the generated code passes all test cases successfully within the first K generations. Following prior studies (Fakhoury et al., 2024; Dong et al., 2024; Jiang et al., 2024), our experimental design adopts K=1, focusing exclusively on first-attempt success rates. This metric also aligns with real-world scenarios where developers aim to produce accurate code on the first attempt (Dong et al., 2024). It should be noted that Pass@1 represents a particularly strict evaluation metric for code generation and improving it is challenging. Higher Pass@1 scores indicate better code generation performance.
3.3. Implementation Detail
3.3.1. Data generation
We use Qwen2.5-Coder-32B-Instruct111https://huggingface.co/Qwen/Qwen2.5-Coder-32B-Instruct as the teacher model with temperature=0.8 to generate incorrect code implementations due to its strong coding abilities and good natural language understanding capabilities. To mitigate the potential threat introduced by errors from the teacher model generation, we manually examined a sample of 50 generated outputs. Our analysis shows that the LLM could produce error-sensitive segments as expected: 96% of the generated incorrect samples are similar to the correct ones, while the remaining 4% completely deviate from correct implementations. Due to space limitations, the checked samples are provided in the replication package. This manual inspection helps validate the reliability of our training data and supports the soundness of our experimental findings.
3.3.2. Settings
All experiments are conducted on a machine with eight Tesla A800 GPUs, each with 80 GB of memory per GPU. is set to 2, which means the weight range of is . All models are trained for 1 epoch. Considering that Fait is designed to enable instruction-tuned LLMs to emphasize error-sensitive segments from their original instruction-tuning datasets, we apply relatively low learning rates during training. Specifically, we use a learning rate of 5e-6 with a linear scheduler and warm-up across all LLMs. The max sequence length is 1024. For inference evaluation, we use greedy decoding to ensure deterministic outputs, which also aligns with prior studies (Chen et al., 2021; Mu et al., 2024).
4. Results
In this section, we report and analyze the experimental results to answer the following research questions (RQs):
RQ1: How effective is our approach in improving code generation across different benchmarks?
RQ2: How do different components of the Fait method contribute to LLMs’ performance?
RQ3: Does Fait demonstrate generalizability across different LLMs and their corresponding instruction-tuning datasets?
RQ4: Does Fait work for instruction-tuned LLMs whose instruction-tuning dataset is closed-source?
4.1. RQ1: Overall Effectiveness
In this RQ, we evaluate the effectiveness of our approach by applying it to several Instruction-tuned LLMs using their corresponding instruction-tuning datasets and assess their performance against three baselines:
Base Models: We use the original instruction-tuned LLMs without any additional Fait as our base models. This comparison demonstrates the absolute improvement achieved through Fait. Specifically, we select three representative instruction-tuned LLMs: MagiCoder-CL (Wei et al., 2024), MagiCoder-DS (Wei et al., 2024) and SemCoder-S (Ding et al., 2024) as our base models.
Closed-Source Models: We include GPT-3.5-Turbo (OpenAI, 2022) and GPT-4-Turbo (OpenAI and Josh Achiam, 2024) as the closed-source baseline. This comparison illustrates the performance gap between our fault-aware fine-tuned LLMs and advanced closed-source LLMs.
Standard-SFT Models: We apply standard SFT on the same base models to create this baseline. This comparison serves two purposes: (1) to examine whether further fine-tuning on coarse-grained instruction-response mappings on their existing dataset can improve performance over the original models, and (2) to highlight the superior performance of our approach in learning fine-grained error-sensitive segments.
We choose the instruction-tuning dataset evol-codealpaca-v1 (Luo et al., 2023) as our training dataset. This dataset is evolved from a seed dataset with GPT-4-Turbo, containing 110K high-quality data points. For MagiCoder-CL and MagiCoder-DS, this dataset is their original instruction-tuning dataset. For SemCoder-S, this dataset is a subset of its original instruction-tuned dataset, which is not fully open-sourced.
Model | Humaneval(+) | MBPP(+) |
Closed-Source Models | ||||||||
GPT-3.5-Turbo (Nov 2023) | 76.8 (70.7) | 82.5 (69.7) | 50.6 | 21.6 | ||||
GPT-4-Turbo (April 2024) | 90.2 (86.6) | 85.7 (73.3) | 58.2 | 35.1 | ||||
Base Model: CodeLlama-Python-7B | ||||||||
MagiCoder-CL | 70.7 (66.5) | 68.4 (56.6) | 39.7 | 12.8 | ||||
+Standard-SFT | 69.5 (64.0) | 69.3 (58.7) | 39.3 | 13.5 | ||||
+Fait | 73.2 (68.9) | 71.7 (59.5) | 42.2 | 15.5 | ||||
Base Model: DeepseekCoder-6.7B-Base | ||||||||
MagiCoder-DS | 76.8 (71.3) | 75.7 (64.4) | 47.6 | 12.8 | ||||
+Standard-SFT | 75.6 (70.7) | 79.1 (66.4) | 46.9 | 10.8 | ||||
+Fait | 77.4 (74.3) | 79.6 (69.0) | 48.2 | 15.5 | ||||
[…]
+Standard-SFT | 79.9 (75.0) | 80.7 (67.2) | 47.1 | 16.2 | ||||
+Fait | 83.5 (78.7) | 83.1 (70.6) | 48.9 | 20.3 |
Table 1 presents the performance of LLMs with Fait and the baselines across HumanEval(+), MBPP(+), and BigCodeBench. Overall, LLMs with Fait demonstrate substantial improvements in code generation. We observe that LLMs with Fait show average relative performance improvements of 4.8% over the base model and 4.9% over LLMs with Standard-SFT on HumanEval(+), MBPP(+), and BigCodeBench. Notably, with Fait, SemCoder-S with only 7B parameters outperforms the closed-source GPT-3.5-Turbo on HumanEval(+) and MBPP(+), and MagiCoder-DS outperforms GPT-3.5-Turbo on HumanEval(+). Both LLMs achieve comparable performance to GPT-3.5-Turbo on BigCodeBench, further validating the exceptional effectiveness of Fait in enhancing code generation capabilities. While the improvement on BigCodeBench-Full is modest, our approach shows more gains on BigCodeBench-Hard (e.g., 20.8% relative improvement for SemCoder-S). This is likely because more challenging problems contain more error-sensitive segments, and Fait is designed to guide LLMs to handle these error-sensitive segments, thus showing greater effectiveness on difficult programming tasks.
When comparing the effects of Fait versus Standard-SFT on base models, we observe that Standard-SFT provides limited improvements and sometimes even weakens the base models. For example, SemCoder-S with Standard-SFT achieves only 0.8% relative performance improvement on HumanEval(+) and suffers 1.9% relative performance decline on MBPP+. This suggests that simply reinforcing the coarse-grained instruction-response mappings on their existing dataset provides minimal benefits, as these models have already captured these general mappings well during their initial instruction tuning. In contrast, Fait goes beyond simply continuing training on the original dataset, enabling LLMs to learn and memorize the mappings between problems and fine-grained error-sensitive segments in code. This targeted approach helps models distinguish between correct implementations and similar-looking but incorrect solutions, thereby generating accurate solutions when encountering target problems.
To further figure out the reasons for Fait improving LLMs’ ability to generate functionally correct code, we manually inspect the test results. Based on our analysis, Fait demonstrates two main advantages over both the original model and Standard-SFT:
First, Fait can learn diverse error-sensitive segments to better recognize and focus on implementation details that are prone to errors, while the original model and Standard-SFT only learn the overall mapping from problem to solution. This method improves the LLM’s attention to key implementation choices. For example, Figure 4 presents a comparison of the results of three versions of SemCoder-S on the HumanEval/135 task. In this example, an error-sensitive segment involves deciding whether to return the index of the target element itself i or the index of its previous element i-1. This distinction directly impacts the functional correctness of the implementation. Both the original model and the model with Standard-SFT incorrectly return the index of the previous element, while the model with Fait correctly returns the index of the target element. We manually examine the tasks from HumanEval(+) that are correctly solved after applying Fait but initially incorrect with the base models. We find that among these tasks, 63.6% of SemCoder-S’s improvements result from properly handling error-sensitive segments, with similar rates observed in MagiCoder-DS (71.4%) and MagiCoder-CL (72.7%). These results further confirm our method’s effectiveness in guiding LLMs to recognize error-sensitive segments.
Second, by developing a deeper understanding of critical code segments, Fait also enhances overall code generation capabilities. By strategically emphasizing error-sensitive segments while maintaining appropriate weight for contextual elements, the LLM learns to identify and handle the crucial parts of implementations that determine correctness. Figure 5 demonstrates this using an example from SemCoder-S on HumanEval/110, which requires determining whether swapping elements between two lists can make all elements in lst1 even. This case highlights improvements that go beyond addressing specific error-sensitive segments. Both the original model and Standard-SFT fail to implement the correct verification logic to determine whether lst2 contains enough even numbers to replace odd numbers in lst1. In contrast, the Fault-Aware Fine-Tuned model correctly implements this logic, demonstrating enhanced general coding abilities rather than just handling error-sensitive parts.
4.2. RQ2: Component Analysis
Model | Humaneval(+) | MBPP(+) |
MagiCoderS-DS | 76.8 (71.3) | 75.7 (64.4) | 47.6 | 12.8 | ||||
+Fait (Line Level) | 75.6 (72.0) | 78.8 (68.3) | 47.1 | 14.2 | ||||
+Fait (Token Level) | 76.2 (72.6) | 79.1 (68.5) | 47.3 | 14.2 | ||||
+Fait | 77.4 (74.3) | 79.6 (69.0) | 48.2 | 15.5 | ||||
SemCoder-S | 79.3 (74.4) | 79.6 (68.5) | 48.5 | 16.9 | ||||
+Fait (Line Level) | 80.5 (76.8) | 83.1 (70.1) | 49.1 | 18.2 | ||||
+Fait (Token Level) | 81.1 (76.8) | 82.8 (70.1) | 48.3 | 16.9 | ||||
+Fait | 83.5 (78.7) | 83.1 (70.6) | 48.9 | 20.3 |
To understand how different components contribute to the effectiveness of Fait, we conduct ablation studies focusing on the impacts of multi-granularity and the loss function in this RQ.
Impact of Difference Granularity. We explore the impact of code difference granularity, which involves synthesizing line-level and token-level code differences to identify error-sensitive segments. Specifically, we conduct ablation experiments using MagiCoder-DS and SemCoder-S as base models and perform evaluation on three selected benchmarks. Table 2 shows the impact of different granularities of differences on Fait. We can observe that the combination of line-level granularity and token-level granularity yields maximum performance gains. For example, when applied to SemCoder-S, this approach achieves a relative average improvement of 4.9% across all benchmarks compared to the base model, compared to just 2.9% for line-level only and 2.4% for token-level only. This demonstrates that multi-granularity differences enable better error-sensitive segments, thereby enhancing LLMs’ code generation capabilities.
Impact of the Loss Function. To validate the effectiveness of our dynamic loss weighting design, we compare our dynamic weighting approach against a fixed weighting strategy where all error-sensitive tokens receive the same constant weight during training. We experiment with fixed weights in the range of on HumanEval(+). Due to the evaluation time constraints, we select SemCoder-S as the representative LLM for this ablation study, as it demonstrates the best overall performance with Fait. Notably, when the fixed weight equals 1, this configuration is equivalent to standard SFT. Figure 6 shows the performance of SemCoder-S with different fixed weights compared to our dynamic weighting approach. Our experimental results reveal two important findings: (1) A fixed weight of 2 yields better performance than other fixed weight values, suggesting that an appropriate constant weight can help the LLM recognize error-sensitive segments and enhance code generation capabilities. (2) Our dynamic weighting approach still outperforms the best fixed weighting configuration. This confirms that dynamically adjusting weights based on the LLM’s current discrimination ability provides better guidance for the LLM to focus on critical implementation details that differentiate correct solutions from their erroneous variants.
4.3. RQ3: The Generalization Capabilities
Model | Humaneval(+) | MBPP(+) |
(Corresponding Dataset OSS-INSTRUCT) | ||||||||
MagiCoder-DS | 66.5 (60.4) | 75.4 (61.9) | 43.4 | 12.2 | ||||
+Standard-SFT | 64.6 (58.5) | 79.1 (66.1) | 43.9 | 12.2 | ||||
+Fait | 67.1 (62.2) | 79.4 (66.4) | 46.2 | 15.5 | ||||
(Corresponding Dataset PYX) | ||||||||
SemCoder | 73.2 (68.9) | 79.9 (65.3) | 43.5 | 16.9 | ||||
+Standard-SFT | 71.3 (65.2) | 79.9 (66.4) | 43.4 | 14.2 | ||||
+Fait | 73.7 (69.5) | 81.0 (67.2) | 47.9 | 21.6 |
In this RQ, we aim to explore the generalizability of Fait across different instruction-tuned LLMs when using their own instruction-tuning datasets. Specifically, we select two representative LLMs and their corresponding instruction datasets. 1) We select MagiCoder-DS and its corresponding dataset OSS-Instruct. This dataset is generated from open-sourced code by GPT-3.5-Turbo, and contains 75K samples. 2) SemCoder and its corresponding dataset PYX. This dataset consists of 95K samples, including comprehensive reasoning texts with executable code samples. The dataset is constructed with problem descriptions generated by GPT-3.5-Turbo and corresponding responses generated by GPT-4o-mini (OpenAI, 2024), creating high-quality instruction-response pairs with detailed reasoning. For each LLM, we process its corresponding dataset through our pipeline and evaluate performance on the same benchmarks used in RQ1. We select Base Models and Standard-SFT Models as our baselines. By using LLMs with different training paradigms and datasets with varying characteristics (open-sourced code versus detailed reasoning with executable samples), we can verify that our approach is not tied to specific LLM series or dataset properties, but rather provides universal benefits.
Table 3 shows the performance of LLMs on Humaneval(+), MBPP(+), and BigCodeBench after Fait and Standard-SFT. We can observe that Fait demonstrates robust generalization capabilities to different instruction-tuning LLMs and their corresponding datasets. For MagiCoder-DS and SemCoder, after Fait, the average performances across all benchmarks show relative improvements of 5.3% and 3.8% compared to the base models. In contrast, standard SFT yielded modest relative improvements of 1.4% for MagiCoder-DS and decreased performance by 2.1% for SemCoder. These results show that Fait’s benefits are not tied to specific dataset characteristics or model series. Instead, the approach effectively enhances diverse instruction-tuned LLMs by teaching them to focus on error-sensitive segments in correct solutions.
4.4. RQ4: Effectiveness on LLMs with Closed-Source Instruction Data
Model | Humaneval(+) | MBPP(+) |
Base Model: CodeLlama-Python-7B | ||||||||
CodeLlama-Instruct | 36.0 (31.1) | 56.1 (46.6) | 25.7 | 4.1 | ||||
+Standard-SFT | 39.0 (34.1) | 61.1 (49.7) | 26.5 | 4.1 | ||||
+Fait | 47.0 (43.9) | 61.9 (51.3) | 29.0 | 4.7 | ||||
Base Model: DeepseekCoder-6.7B-Base | ||||||||
DeepseekCoder-Instruct | 73.8 (70.7) | 74.9 (65.6) | 43.8 | 15.5 | ||||
+Standard-SFT | 75.6 (70.1) | 77.8 (66.9) | 42.0 | 13.5 | ||||
+Fait | 81.7 (76.2) | 78.6 (66.9) | 44.3 | 16.9 |
A key question for the broader adoption is whether Fait can enhance LLMs whose original instruction-tuning datasets are not publicly available. To investigate this, we applied our method to two widely-used LLMs with closed-source training data in this RQ. Specifically, we choose CodeLlama-7B-Instruct (Roziere et al., 2023) and DeepSeek-
Coder-6.7B-Instruct (Guo et al., 2024) as base models. These LLMs are instruction-tuned on substantial but proprietary datasets - CodeLlama-7B-Instruct underwent instruction tuning on approximately 5B tokens of instruction data, while DeepSeekCoder-6.7B-Instruct is tuned on around 2B tokens. To test our approach without access to these original datasets, we select evol-codealpaca-v1 used in the main experiment as the training dataset. We select Base Models and Base Models with Standard SFT as our baselines and evaluate on Humaneval(+), MBPP(+) and BigCodeBench(+).
Table 4 shows the performance of these two LLMs using Standard-SFT and fault-fine-tuning on Humaneval(+), MBPP(+), and BigCodeBench. We can find that Fait is also applicable to LLMs with closed-source datasets. For CodeLlama-Instruct and DeepseekCoder-Instruct, after Fait, the average relative improvements across all benchmarks are 19.1% and 5.9%. By comparison, the standard SFT yield gains of 7.5% and 0.5%, respectively. This further demonstrates that fault-fine-tuning is also applicable to LLMs with closed-source datasets, showing strong applicability.
5. Discussion
Impact of in Loss. The parameter in our loss function controls the emphasis placed on error-sensitive tokens. We further experimentally investigate the performance impact of changing . Specifically, we conduct experiments with on Humaneval(+) and observe the performance changes of LLMs. Due to the evaluation time constraints, we select SemCoder-S as the representative LLM for this ablation study, as it demonstrates the best overall performance with Fait. Figure 7 illustrates the performance trends as varies. We can observe that provides an optimal balance between emphasizing error-sensitive tokens and maintaining attention to shared tokens. Additionally, it can be observed that in all cases, after Fait, the LLM’s performance matches or exceeds that of Standard-SFT, demonstrating the robustness to hyperparameter choices.
Model | Humaneval(+) | MBPP(+) |
Base Model: CodeLlama-Python-7B | ||||||||
MagiCoder-CL | 70.7 (66.5) | 68.4 (56.6) | 39.7 | 12.8 | ||||
+DPO | 66.5 (61.6) | 68.8 (58.7) | 39.8 | 14.2 | ||||
+Fait | 73.2 (68.9) | 71.7 (59.5) | 42.2 | 15.5 | ||||
Base Model: DeepseekCoder-6.7B-Base | ||||||||
MagiCoder-DS | 76.8 (71.3) | 75.7 (64.4) | 47.6 | 12.8 | ||||
+DPO | 76.2 (71.9) | 79.1 (68.3) | 47.8 | 13.5 | ||||
+Fait | 77.4 (74.3) | 79.6 (69.0) | 48.2 | 15.5 | ||||
[…]
+DPO | 81.7 (76.2) | 81.0 (67.7) | 47.9 | 16.2 | ||||
+Fait | 82.9 (79.3) | 83.1 (70.6) | 48.9 | 20.3 |
Compared to Reinforcement Learning Method. Given the increasing popularity of reinforcement learning (RL) methods for improving code generation (Le et al., 2022; Wang et al., 2022), we believe it’s important to compare our approach with these established techniques. As RL methods in code generation typically aim to align model outputs with desired code solutions by increasing the probability of correct implementations while reducing the likelihood of erroneous ones, they share similarities with our work principle of enhancing LLMs’ ability to identify error-sensitive segments to improve code generation capabilities. Specifically, we compare our approach with the representative DPO method (Rafailov et al., 2023), which is widely used and has demonstrated significant advantages in code generation (Yang et al., 2024; Miao et al., 2024; Gee et al., 2024; Zhang et al., 2024; Gallego, 2024). This method works by training models to directly maximize the likelihood of preferred outputs over non-preferred ones without requiring explicit reward modeling, learning from paired examples of more and less desirable code implementations.
To ensure a fair comparison, we select the same LLMs and use identical experimental settings as stated in RQ1 for DPO training. The training dataset remains consistent across both DPO and Fait, and we evaluate and compare the performance of DPO and Fait on the Humaneval(+), MBPP(+), and BigCodeBench benchmarks.
Table 5 shows the performance of LLMs trained with different methods. Overall, LLMs with Fait consistently outperform those trained with DPO. We can observe that Fait outperforms DPO by a relative average of 4.2%, across three selected benchmarks. This advantage stems from fundamental methodological differences: while DPO relies on coarse-grained preference signals that cannot precisely target error-sensitive segments, Fait specifically maintains learning across all tokens while strategically emphasizing error-sensitive segments within code implementations. This approach ensures the LLM retains general coding knowledge while becoming more attentive to critical details that often determine functional correctness.
Additionally, we note that DPO’s ability to differentially reward correct implementations and penalize incorrect ones could be leveraged to enhance the learning of error-sensitive segments. Specifically, a tailored reward function could be designed to strengthen the model’s focus on these critical segments, potentially combining the strengths of both approaches. We leave this promising direction for future exploration.
6. Threats To Validity
There are three major threats to the validity of our work.
Threats to external validity relate to the generalizability of our approach. While we evaluate our approach on multiple instruction-tuned models, there may be concerns about generalization to other LLMs. However, this threat is mitigated by our diverse selection of models with different series. Furthermore, the cross-dataset experiments (RQ3) and closed-source dataset experiments (RQ4) demonstrate robust generalization capabilities across different settings. In addition, due to computational resource constraints, our experiments primarily focus on 7B parameter LLMs rather than larger LLMs. In future work, we plan to explore a broader range of model series to further validate our approach’s generalizability.
Threats to internal validity involve the impact of the quality of incorrect code and choices of hyperparameters. The effectiveness of our approach depends on the quality of incorrect code variants and the weighting factor . To mitigate threats related to code quality, we prompt a strong teacher model (Qwen2.5-Coder-32B-Instruct) to generate plausible incorrect variants and manually verify whether the generated data meets our expectations. While a small portion of noise data - incorrect implementations that differ completely from correct solutions - remains present, we argue these instances may actually enhance model robustness by preventing overfitting to specific error-sensitive segments (DeVries and Taylor, 2017; Xie et al., 2020). For hyperparameter-related threats, we conduct extensive sensitivity analysis as shown in Figure 7. In future work, we intend to investigate the use of stronger teacher models, such as GPT-4-Turbo, to generate similar incorrect code and examine their impact.
Threats to construct validity relate to the reliability of evaluation metrics. We evaluate our approach using the pass rates metric; however, this metric may inadequately capture the functional correctness of generated code implementations with a limited number of test cases. To address this limitation, we deliberately incorporate the extended versions of some benchmarks, which substantially expand the number of test cases. In future work, we plan to explore additional evaluation approaches, such as LLM-as-a-Judge (Gu et al., 2024), to provide a more comprehensive assessment of code quality beyond functional correctness.
7. Related Work
7.1. LLMs for Code Generation
As a momentous milestone, Codex (Chen et al., 2021) boasting a 12-billion-parameter model demonstrates the extraordinary capability to tackle up to 72% of Python programming problems. After that, a new wave of code generation models, such as AlphaCode (Li et al., 2022), CodeGen (Nijkamp et al., [n. d.]), InCoder (Fried et al., [n. d.]) and StarCoder (Li et al., 2023) are proposed and have shown promising results in the code generation task. Building upon these foundations, more code-focused LLMs emerged, such as Magicoder (Wei et al., 2024), SemCoder (Ding et al., 2024), WaveCoder (Yu et al., 2024) and WizardCoder (Luo et al., [n. d.]). These specialized LLMs are typically based on general LLMs in solving domain-specific coding tasks through instruction tuning.
7.2. Fine-tuning on Code LLM
Fine-tuning pre-trained language models has emerged as a dominant paradigm for optimizing performance in code generation. Instruction tuning (Ouyang et al., 2022; Achiam et al., 2023), as a form of supervised fine-tuning, aims to align LLMs with instruction through high-quality instruction corpora. For instance, Magicoder (Wei et al., 2024) introduce OSS-Instruct, an instruction tuning dataset generated by a teacher LLM drawing inspiration from open-source code snippets, which effectively enhances code generation capabilities when used for fine-tuning. Furthermore, OSS-Instruct is orthogonal to existing instruction tuning datasets like Evol-Instruct (Luo et al., [n. d.]), enabling the Magicoder series LLMs finetuned on this combined data to achieve further performance improvements. Similarly, SemCoder (Ding et al., 2024) propose PYX, a dataset created by a teacher LLM simulating human debugging processes. By incorporating data that simulates execution reasoning and captures code execution nuances, LLMs finetuned with PYX understand and articulate the execution process step-by-step, enhancing their reasoning capabilities. Likewise, combining with existing instruction tuning datasets like Evol-Instruct, the resulting SemCoder-S LLM further improves the original LLM’s code generation abilities.
To address limitations in preventing untruthful and unexpected outputs, researchers explore reinforcement learning (Ouyang et al., 2022). To address limitations in undesired outputs, researchers have explored reinforcement learning approaches. For instance, CodeRL (Le et al., 2022) utilizes compiler feedback as reward signals combined with REINFORCE (Williams, 1992) to fine-tune CodeT5 (Wang et al., 2021), reducing compilation errors. SENSE (Yang et al., 2024) synthesizes text-to-SQL training data from both strong and weak language models, and leverage Direct Preference Optimization (DPO) (Rafailov et al., 2023) to learn from correct and incorrect SQL examples, demonstrating state-of-the-art performance on text-to-SQL benchmarks. Similarly, PPOCoder (Shojaee et al., [n. d.]) trains CodeT5 with proximal policy optimization (Schulman et al., 2017). However, despite these advances, these fine-tuning approaches face a fundamental limitation: they treat all tokens with equal importance during loss calculation, making it difficult for models to distinguish semantically correct implementations from syntactically similar but incorrect ones. In this paper, we aim to address this challenge in code LLMs.
We find recent work Focused-DPO (Zhang et al., 2025) enhances code generation capability by concentrating preference optimization on error-prone code points through an improved DPO (Rafailov et al., 2023) methodology. Our method is complementary to this approach - while Focused-DPO operates during post-training reinforcement learning stages, our approach operates during the supervised fine-tuning stage. Additionally, their work has demonstrated that their method can strengthen the capabilities of post-trained models. However, as Focused-DPO is currently under peer review and its implementation is not yet publicly available, we are unable to experimentally validate the potential synergies between our approaches in this work.
8. Conclusion
In this paper, we introduce Fault-Aware Fine-Tuning (Fait), a novel fine-tuning technique that enhances code generation capabilities in instruction-tuned LLMs by refining their ability to distinguish between correct implementations and subtly incorrect variants. Unlike conventional supervised fine-tuning that treats all tokens equally, our approach identifies and prioritizes error-sensitive segments through two key components: error-sensitive segments identification, that captures both line-level and token-level critical differences, and dynamic importance reweighting that dynamically adjusts token weights during training to focus on discriminative elements. Through extensive experiments across seven LLMs and three widely-used benchmarks, we demonstrate that our method achieves an average relative improvement of 6.9% on pass@1 with just one epoch of training, with certain enhanced 6.7B LLMs even outperforming GPT-3.5-Turbo on selected benchmarks. The technique also exhibits strong generalization capabilities across diverse instruction-tuned LLMs and maintains effectiveness even when applied to LLMs with closed-source instruction datasets, providing a practical solution to improve LLMs’ code generation capabilities.
Data Availability
Our code is available: https://anonymous.4open.science/r/FAFT-976B.